OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)? Possible answers: -x + y = 2 -x – y= 2 x - y = 2 -x - y = 0 x + y = 0

OpenStudy (anonymous):

the slope of a perpendicular line is the negative reciprocal of the slope of the given line -x + y =7, so the slope is 1. Negative reciprocal of this is -1. therefore, you have x+y=b solve your b using your point. you'll get: x + y =0

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

np :)